OpenStudy (anonymous):

y=3/(squareroot3x+4) Find f(x) and g(x) so that the function can be described as y = f(g(x)).

OpenStudy (anonymous):

you have lots and lots of choices

OpenStudy (anonymous):

The choices are: f(x) = 3/squarerootx, g(x) = 3x + 4 f(x) = squareroot3x+4, g(x) = 3 f(x) = 3/x, g(x) = 3x + 4 f(x) = 3, g(x) = squareroot3x+4

OpenStudy (anonymous):

Sorry. I don't understand how to know what to extract/place to find f((g)x))

OpenStudy (anonymous):

the first one is one choice for sure

OpenStudy (anonymous):

if \(f(x)=\frac{3}{\sqrt{x}}\) and \(g(x)=3x+4\) then \[f\circ g(x)=f(g(x))=f(3x+4)=\frac{3}{\sqrt{3x+4}}\]

OpenStudy (anonymous):

but it is silly to give you choices because you have lots of other options of those choices however, that is the only one that works

OpenStudy (anonymous):

Oh I understand what you did! Thanks so much!

OpenStudy (anonymous):

yw